# Aligning Model Properties via Conformal Risk Control

### v1

@@ Results.
Table 1 presents results on the test set for the Combined Cycle Power Plant dataset (Tfekci and Kaya, 2014). In practice, Exhaust-vacuum is known to negatively influence turbine efficiency (Tfekci and Kaya, 2014). The conformal procedure outperforms the constrained model in terms of MSE for $\alpha=0.1$ and $\alpha=0.05$, but worsens for $\alpha=0.01$. The risk metric closely matches the theoretical guarantee from conformal risk control and achieves optimal performance of 0 for the constrained model. Additional datasets and results are detailed in the appendix.
Table 1: Power Plant, $n=9568$. Monotonically decreasing on Exhaust Vacuum. $\boldsymbol{\lambda}^{\max}=(10,10)$, $g(\boldsymbol{\lambda})=\lambda^{+}+\lambda^{-}$.
$\alpha$ $\boldsymbol{\lambda}$ Metric Unconstrained Conformal Constrained
0.1 $\lambda^{+}=0.96$ MSE 9.99 10.35 15.31
$\lambda^{-}=1.18$ Risk 0.17 0.111 0
0.05 $\lambda^{+}=1.94$ MSE 9.99 11.37 15.31
$\lambda^{-}=2.68$ Risk 0.17 0.058 0
0.01 $\lambda^{+}=4.76$ MSE 9.99 15.56 15.31
$\lambda^{-}=4.88$ Risk 0.17 0.012 0
Figure 1: Univariate partial dependence plot of unconstrained model. Risk control band for $\alpha=0.05$. Dashed line exemplifying Theorem 1 demonstrating existence of monotonically decreasing function falling within the conformal band on $0.97>1-\alpha$ fraction of the domain.
@@ 4.2 Concavity
Concavity and convexity are crucial behaviors in many applications. In this context, we focus on concavity in a single feature. A common example where users might expect concave behavior is in recommendation or preference prediction models. According to economic theory, the utility function with respect to the quantity of an item is often quasi-concave, reflecting the principle of diminishing marginal utility (Mas-Colell et al., 1995). Jenkins et al. (2021) propose a novel loss function to account for this expected concavity, which aligns the model with the concavity property ${\mathcal{P}}$ during training. Here we again consider aligning a pre-trained model, not trained to satisfy ${\mathcal{P}}$, using a proximity oblivious tester $\mathsf{T}$ for ${\mathcal{P}}$ as described in Algorithm 2.
Algorithm 2 POT $\mathsf{T}$ for concavity in a fixed dimension $k$
1:  Sample $(X,Y)\sim\mathcal{D}$. Let $X=(x,x^{-k})$
2:  Let $L$ and $R$ be the subset of points in dimension $k$ less than $x$ and greater than $x$, respectively.
3:  Sample $x_{l}$ uniformly from $L$ and $x_{r}$ uniformly from $R$.
4:  Set $X_{l}=(x_{l},x^{-k}),X_{r}=(x_{r},x^{-k})$
5:  if $f(X)<\frac{f(X_{l})+f(X_{r})}{2}$ then
6:     return  Reject
7:  end if
8:  return  Accept
In this case we use a scalar $\lambda$ and have that that $L=\ell(C_{\lambda}(X),Y)=0$ if $f(X)+\lambda\geq\frac{f(X_{l})+f(X_{r})}{2}$, and $L=\ell(C_{\lambda}(X),Y)=1$ otherwise. We demonstrate running conformal risk control with this loss function on a real-world dataset in the appendix.
@@ 5 A stylized examination of alignment persistence in AI models
In Appendix C, we utilize results from Misiakiewicz and Montanari (2023) to derive insights into the impact of model complexity and data size on adherence to ${\mathcal{P}}$. In particular, we show that under certain assumptions, including small noise bias and robustness of property ${\mathcal{P}}$, the constrained and unconstrained models have zero distance in the classic setting: $\hat{{\mathbf{a}}}_{\lambda,{\mathcal{P}}}=\hat{{\mathbf{a}}}_{\lambda}$. However, in the underspecified setting, the constrained and unconstrained models will differ, resulting in a non-zero distance: $\hat{{\mathbf{a}}}_{\lambda,{\mathcal{P}}}\neq\hat{{\mathbf{a}}}_{\lambda}$. This result implies that in the presence of noise bias, the overparameterized models (i.e., underspecified setting) fail to satisfy the property ${\mathcal{P}}$, and this cannot be remedied as the data size increases.
@@ 6 Related work
Connections to Yadkori et al. (2024).
The work of Yadkori et al. (2024) closely aligns with ours in both methodology and theme, utilizing conformal risk control to reduce LLM hallucinations (Ji et al., 2023). To demonstrate our framework’s flexibility and generality, we apply it to their context. They examine a potentially random LLM $f:{\mathcal{X}}\to\mathcal{Y}$ and assume access to a confidence score function $g:{\mathcal{X}}\to\mathbb{R}$. They use an abstention function $a:\Lambda\times{\mathcal{X}}\to\{0,1\}$ that decides whether the model should abstain from answering where $a_{\lambda}(X)=1$ if $g(X)<\lambda$ and $a_{\lambda}(X)=0$ if $g(X)\geq\lambda$ The match function $m:{\mathcal{X}}\times\mathcal{Y}\times\mathcal{Y}\to\{0,1\}$ is used to determine if a response $Y^{\prime}$ is semantically equivalent to $Y$, indicating hallucination if not.
Yadkori et al. (2024) use conformal risk control to find an optimal $\hat{\lambda}$ such that the pair $(a_{\hat{\lambda}},f)$ hallucinates at a rate less than $\alpha$. The loss function they use is $\ell:{\mathcal{X}}\times\mathcal{Y}\times\Lambda\to\mathbb{R}$ such that $\ell$ punishes failure to abstain when the response does not match the label, that is
$\ell(X,Y;\lambda)=(1-a_{\lambda}(X))(1-m(X;f(X),Y)).$ (1)
Yadkori et al. (2024) use a separate conformal procedures for setting $\lambda$ and a parameter $\beta$ with which they define $m(X;Y^{\prime},Y)=\mathbb{I}(s(X;Y^{\prime},Y)\geq\beta)$, where the similarity function $s:{\mathcal{X}}\times{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is assumed to be given. We can use our multi-lambda conformal risk control with $\boldsymbol{\lambda}=(\lambda_{a},\lambda_{m})$ to set these parameters simultaneously using a single calibration set $\{(X_{i},Y_{i})\}_{i=1}^{n}$ of ground truth query-response pairs sampled from the true data distribution ${\mathcal{D}}$. Let
@@ 7 Discussion
We introduce a method to align pre-trained models with desired user properties using conformal risk control. By post-processing outputs using property dependent loss functions, we provide probabilistic guarantees that conformal intervals contain functions close to the desired set. This allows for alignment without retraining, effective in both generative and non-generative contexts. Future work should extend these techniques to more properties, explore sample complexity and adaptive querying, and potentially apply them to policy functions in MPD settings for RL agent safety guarantees.
Appendix A Proof of multi-lambda conformal risk control

### v2

@@ Results.
Table 1 presents results on the test set for the Combined Cycle Power Plant dataset (Tfekci and Kaya, 2014). In practice, Exhaust-vacuum is known to negatively influence turbine efficiency (Tfekci and Kaya, 2014). The conformal procedure outperforms the constrained model in terms of MSE for all $\alpha$, which is a fortuitous but unexpected outcome. The constrained model should be seen as an oracle benchmark in the sense that the model was given to the user already trained to satisfy the desired property. The results on this dataset and The risk metric closely matches the theoretical guarantee from conformal risk control and achieves optimal performance of 0 for the constrained model. Additional datasets and results are detailed in the appendix.
Table 1: Power Plant, $n=9568$. Monotonically decreasing on Exhaust Vacuum. $\boldsymbol{\lambda}^{\max}=(10,10)$.
$\alpha$ $\boldsymbol{\lambda}$ Metric Unconstrained Adjusted Constrained
$\lambda^{+}=0.51_{\scriptsize{(\pm 0.24)}}$ MSE $10.19_{\scriptsize{(\pm 0.46)}}$ $10.47_{\scriptsize{(\pm 0.46)}}$ $16.21_{\scriptsize{(\pm 0.45)}}$
0.1 $\lambda^{-}=0.76_{\scriptsize{(\pm 0.24)}}$ Risk $0.75_{\scriptsize{(\pm 0.09)}}$ $0.10_{\scriptsize{(\pm 0.001)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
$\lambda^{+}=1.09_{\scriptsize{(\pm 0.51)}}$ MSE $10.19_{\scriptsize{(\pm 0.46)}}$ $11.42_{\scriptsize{(\pm 0.44)}}$ $16.21_{\scriptsize{(\pm 0.45)}}$
0.05 $\lambda^{-}=1.61_{\scriptsize{(\pm 0.50)}}$ Risk $0.75_{\scriptsize{(\pm 0.09)}}$ $0.05_{\scriptsize{(\pm 0.001)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
$\lambda^{+}=2.39_{\scriptsize{(\pm 0.82)}}$ MSE $10.19_{\scriptsize{(\pm 0.46)}}$ $14.46_{\scriptsize{(\pm 0.48)}}$ $16.21_{\scriptsize{(\pm 0.45)}}$
0.01 $\lambda^{-}=3.33_{\scriptsize{(\pm 0.79)}}$ Risk $0.75_{\scriptsize{(\pm 0.09)}}$ $0.01_{\scriptsize{(\pm 0.001)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
Figure 1: Univariate partial dependence plot of unconstrained model. Risk control band for $\alpha=0.05$. Dashed line exemplifying Theorem 1 demonstrating existence of monotonically decreasing function falling within the conformal band on $0.975>1-\alpha$ fraction of the domain.
@@ 4.2 Concavity
Concavity and convexity are crucial behaviors in many applications. In this context, we focus on concavity in a single feature. A common example where users might expect concave behavior is in recommendation or preference prediction models. According to economic theory, the utility function with respect to the quantity of an item is often quasi-concave, reflecting the principle of diminishing marginal utility (Mas-Colell et al., 1995). Jenkins et al. (2021) propose a novel loss function to account for this expected concavity, which aligns the model with the concavity property ${\mathcal{P}}$ during training. Here we again consider aligning a pre-trained model, not trained to satisfy ${\mathcal{P}}$, using a proximity oblivious tester ${\mathcal{T}}$ for ${\mathcal{P}}$ as described in Algorithm 2.
Algorithm 2 POT $\mathcal{T}$ for property $\mathcal{P}$ of concavity in dimension $k$
1:  Sample $X_{\text{mid}}\sim\mathcal{D}$
2:  Sample $\delta_{\text{left}},\delta_{\text{right}}$ from empirical differences in feature $k$
3:  Set $X_{\text{left}}$ by decreasing feature $k$ of $X_{\text{mid}}$ by $\delta_{\text{left}}$
4:  Set $X_{\text{right}}$ by increasing feature $k$ of $X_{\text{mid}}$ by $\delta_{\text{right}}$
5:  Query $F(X_{\text{mid}})$, $F(X_{\text{left}})$, and $F(X_{\text{right}})$
6:  Compute $\alpha=\dfrac{X_{\text{right}}[k]-X_{\text{mid}}[k]}{X_{\text{right}}[k]-X_{\text{left}}[k]}$
7:  if  $\min F(X_{\text{mid}})>\alpha\max F(X_{\text{left}})+(1-\alpha)\max F(X_{\text{right}})$  then
8:     return  Reject
9:  end if
10:  return  Accept
[…]
@@ 5 A stylized examination of alignment persistence in AI models
In Appendix E, we utilize results from Misiakiewicz and Montanari (2023) to derive insights into the impact of model complexity and data size on adherence to ${\mathcal{P}}$. In particular, we show that under certain assumptions, including small noise bias and robustness of property ${\mathcal{P}}$, the constrained and unconstrained models have zero distance in the classic setting: $\hat{{\mathbf{a}}}_{\lambda,{\mathcal{P}}}=\hat{{\mathbf{a}}}_{\lambda}$. However, in the underspecified setting, the constrained and unconstrained models will differ, resulting in a non-zero distance: $\hat{{\mathbf{a}}}_{\lambda,{\mathcal{P}}}\neq\hat{{\mathbf{a}}}_{\lambda}$. This result implies that in the presence of noise bias, the overparameterized models (i.e., underspecified setting) fail to satisfy the property ${\mathcal{P}}$, and this cannot be remedied as the data size increases.
@@ 6 Related work
The work of Yadkori et al. (2024) closely aligns with ours in both methodology and theme, utilizing conformal risk control to reduce LLM hallucinations (Ji et al., 2023). We discuss connections to this work in Appendix  F.
@@ 7 Discussion
We introduce a method to align pre-trained models with desired user properties using conformal risk control. By post-processing outputs using property dependent loss functions, we provide probabilistic guarantees that conformal intervals contain functions close to the desired set. This allows for alignment without retraining, effective in both generative and non-generative contexts. Future work should extend these techniques to more properties, explore sample complexity and adaptive querying, and potentially apply them to policy functions in MDP settings for RL agent safety guarantees.
Appendix A Property Testing
In this section, we provide a brief introduction to property testing, drawing upon standard definitions and concepts from the field. Property testing is a framework for designing algorithms that quickly decide whether a large, complex object (such as a function or a graph) possesses a certain property or is far from having that property. The goal is to make this decision by inspecting only a small, random portion of the object, thereby significantly reducing the computational effort required compared to examining the entire object.
Our definitions and explanations are based on Goldreich [2017], which offers a comprehensive introduction to property testing. We focus on the basic notions essential for understanding property testing as used in this paper, including the concepts of distance between objects, testers, and proximity-oblivious testers.
A.1 Distance and $\varepsilon$-Farness
A central concept in property testing is the notion of distance between objects, which allows us to formalize what it means for an object to be far from satisfying a property. For functions, this distance is typically measured with respect to a distribution over the input domain.
Definition 7 ($\varepsilon$-far).
Let $f:\mathcal{X}\to\mathcal{Y}$ be a function, $\mathcal{D}$ a distribution over $\mathcal{X}$, $\varepsilon>0$, and $\mathcal{P}$ a property (a set of functions). We say that $f$ is $\varepsilon$-far from $\mathcal{P}$ with respect to $\mathcal{D}$ if the distance $\delta_{\mathcal{P},\mathcal{D}}(f)>\varepsilon$, where
$\delta_{\mathcal{P},\mathcal{D}}(f)=\inf_{g\in\mathcal{P}}\delta_{\mathcal{D}}(f,g),\quad\text{and}\quad\delta_{\mathcal{D}}(f,g)=\Pr_{X\sim\mathcal{D}}[f(X)\neq g(X)].$
In other words, a function $f$ is $\varepsilon$-far from $\mathcal{P}$ if any function $g\in\mathcal{P}$ agrees with $f$ on at most a $1-\varepsilon$ fraction of inputs sampled from $\mathcal{D}$. This notion of distance allows us to quantify how much $f$ differs from satisfying the property $\mathcal{P}$.
While many property testing algorithms assume $\mathcal{D}$ to be the uniform distribution over $\mathcal{X}$, in some settings, particularly when dealing with arbitrary data distributions, it is natural to consider general distributions. This leads to the distribution-free property testing model Goldreich [2017], Halevy and Kushilevitz [2007], where algorithms are designed to work with any underlying distribution $\mathcal{D}$.
A.2 Testers and Error Types
Property testing algorithms, or testers, aim to distinguish between objects that have a certain property and those that are $\varepsilon$-far from it, by querying the object at a small number of locations. Testers can be categorized based on their error probabilities:
  • One-sided error testers: These testers always accept objects that have the property (zero false negatives), but only reject objects that do not satisfy the property with some probability that may be below 1.
  • Two-sided error testers: These testers may err in both directions, accepting objects that are $\varepsilon$-far from the property or rejecting objects that have the property, but with bounded error probabilities.
Technically such testers are referred to as oracle machines. An oracle machine is a theoretical computational model, specifically a Turing machine that has access to an oracle. The oracle is a black box that can compute certain functions or provide certain information that the machine cannot compute on its own. In our context, the oracle machine $\mathcal{M}$ can query the function $F$ at any point $x\in\mathcal{X}$ to receive $F(x)$ and can obtain samples from the distribution $\mathcal{D}$. It does not have explicit knowledge of the internal structure of $F$ or $\mathcal{D}$ but can interact with them through these oracle queries.
For our purposes in this work we will only need to discuss one-sided error testers, which we can formally define as follows:
Definition 8 (One-Sided Error Tester).
A one-sided error tester for a property $\mathcal{P}$ is a probabilistic oracle machine $\mathcal{M}$ that, given a distance parameter $\varepsilon>0$, oracle access to a function $f:\mathcal{X}\to\mathcal{Y}$, and oracle access to samples from a distribution $\mathcal{D}$ over $\mathcal{X}$, satisfies:
  1. 1.
    If $f\in\mathcal{P}$, then $\Pr[\mathcal{M}^{f,\mathcal{D}}(\varepsilon)=\text{Accept}]=1$.
  2. 2.
    If $f$ is $\varepsilon$-far from $\mathcal{P}$ with respect to $\mathcal{D}$, then $\Pr[\mathcal{M}^{f,\mathcal{D}}(\varepsilon)=\text{Accept}]\leq\frac{1}{3}$.
The choice of error probability $\frac{1}{3}$ is arbitrary and can be reduced to any $\delta\in(0,\frac{1}{2})$ by repeating the tester multiple times and taking the majority outcome.
A.3 Proximity-Oblivious Testers
In many property testing algorithms, the proximity parameter $\varepsilon$ is used only to determine the number of iterations or samples needed, rather than affecting the core logic of the tester. This observation leads to the concept of proximity-oblivious testers (POTs), where the basic testing procedure is independent of $\varepsilon$.
Definition 9 (Proximity-Oblivious Tester).
A proximity-oblivious tester for a property $\mathcal{P}$ is a probabilistic oracle machine $\mathcal{T}$ that operates without knowledge of the proximity parameter $\varepsilon$ and satisfies:
  1. 1.
    If $f\in\mathcal{P}$, then $\Pr[\mathcal{T}^{f,\mathcal{D}}=\text{Accept}]=1$
  2. 2.
    If $f\notin\mathcal{P}$, then $\Pr[\mathcal{T}^{f,\mathcal{D}}=\text{Reject}]\geq\rho(\delta_{\mathcal{P},\mathcal{D}}(f))$, where $\rho:(0,1]\to(0,1]$ is a non-decreasing function called the detection probability.
Note that the probability of rejecting a function increases with its distance from $\mathcal{P}$, as quantified by the detection probability function $\rho$.
To convert a POT into a standard tester with error probability $\frac{1}{3}$, we can repeat the basic testing procedure a sufficient number of times, determined by $\rho(\varepsilon)$, and accept or reject based on the aggregate outcomes. Making $O\left(\frac{1}{\rho(\varepsilon)}\right)$ independent calls to $\mathcal{T}$ and accepting if all calls accept yields a tester that satisfies the standard definition with the desired error probability Goldreich and Ron [2008].
A.4 Adaptivity and Non-Adaptivity
Testers can also be classified based on whether they are adaptive or non-adaptive in their querying strategy:
  • Non-adaptive testers: The queries made by the tester are determined in advance, based solely on the random coins and explicit inputs, and do not depend on the answers to previous queries.
  • Adaptive testers: The tester’s queries may depend on the answers to previous queries, allowing for potentially more powerful testing strategies.
Non-adaptive testers are often simpler and easier to analyze, and many property testing algorithms are designed to be non-adaptive. In the context of POTs, non-adaptivity is particularly advantageous because the tester’s behavior remains consistent regardless of the function being tested. In this work we focus only on non-adaptive testers, but exploring the use of adaptive testers is an exciting future direction.
Appendix B Proofs for Proximity Oblivious Testers
We provide proofs that each of the testing algorithms we use are indeed Proximity Oblivious Testers according to Definition 4. However we note that these properties are well studied in the $f:{\mathcal{X}}\to{\mathcal{Y}}$ case and thus we provide these proofs in the set-valued $f:{\mathcal{X}}\to 2^{{\mathcal{Y}}}$ case mainly for completeness.
B.1 Proximity Oblivious Tester for constant property
For the sake of convenience we repeat the setting from the Example in Section 2.1.
Consider functions $f:\mathbb{R}^{d}\to\mathbb{R}$, and let $\mathcal{P}$ denote the property that $f$ is constant in the $k$-th dimension. Assume ${\mathcal{D}}$ is the empirical distribution of the inputs $X\in\mathbb{R}^{d}$ for some fixed dataset.
Restrict to set-valued functions $F$ that output compact and connected intervals of the form $[a,b]\subseteq\mathbb{R}$ for $a,b\in\mathbb{R}$. The candidate POT $\mathcal{T}^{F,\mathcal{D}}$ for whether such a set-valued function $F$ accommodates ${\mathcal{P}}$ is then as follows: sample $X,X^{\prime}\sim\mathcal{D}$, If $F(X)\cap F(X^{\prime})\neq\varnothing$, then Accept; otherwise, Reject. We claim that this is indeed a POT.
Proof.
Assume that $F$ accommodates $\mathcal{P}$. This means there exists a function $g\in\mathcal{P}$ such that for all $X\in\mathcal{X}$, $g(X)\in F(X).$ Since $g$ is constant in the $k$-th dimension, there exists a function $h:\mathbb{R}^{d-1}\to\mathbb{R}$ such that for all $x_{k}\in\mathbb{R}$ and $x^{-k}\in\mathbb{R}^{d-1}$, $g(x_{k},x^{-k})=h(x^{-k}).$ And in the algorithm we sample $X=(x_{k},x^{-k})$ from $\mathcal{D}$ and we independently sample $x_{k}^{\prime}$ from the marginal distribution of $x_{k}$ and set $X^{\prime}=(x_{k}^{\prime},x^{-k})$. Then since $x^{-k}$ is the same for both $X$ and $X^{\prime}$, and $g$ is constant in $x_{k}$, we have:
$g(X)=h(x^{-k})=g(X^{\prime}).$
Because $g(X)\in F(X)$ and $g(X^{\prime})\in F(X^{\prime})$, it follows that:
$h(x^{-k})\in F(X)\cap F(X^{\prime}).$
Therefore, $F(X)\cap F(X^{\prime})\neq\emptyset$, and the algorithm returns Accept.
Now suppose that $F$ is $\varepsilon$-far from $\mathcal{P}$. This means that for any function $g\in\mathcal{P}$,
$\delta_{\mathcal{D}}(F,g)=\Pr_{X\sim\mathcal{D}}[g(X)\notin F(X)]>\varepsilon.$
For each $g\in\mathcal{P}$, define the set of points where $g$ does not belong to $F(X)$:
$S_{g}=\{X\in\mathcal{X}:g(X)\notin F(X)\}.$
Although the sets $S_{g}$ may vary with $g$, for each $g\in\mathcal{P}$, we have $\mathcal{D}(S_{g})>\varepsilon.$ Fix an arbitrary function $g\in\mathcal{P}$. For a fixed $x^{-k}\in\mathbb{R}^{d-1}$, define:
$p_{g}(x^{-k})=\Pr_{x_{k}\sim\mathcal{D}_{x_{k}|x^{-k}}}[g(x_{k},x^{-k})\notin F(x_{k},x^{-k})],$
where $\mathcal{D}_{x_{k}|x^{-k}}$ is the conditional distribution of $x_{k}$ given $x^{-k}$.
Since $\delta_{\mathcal{D}}(F,g)=\mathbb{E}_{x^{-k}}[p_{g}(x^{-k})]>\varepsilon$, it follows that the average error probability over $x^{-k}$ is greater than $\varepsilon$.
When we sample $X=(x_{k},x^{-k})$ and $X^{\prime}=(x_{k}^{\prime},x^{-k})$ with the same $x^{-k}$ and independent $x_{k}$ and $x_{k}^{\prime}$, the probability that both $g(X)\notin F(X)$ and $g(X^{\prime})\notin F(X^{\prime})$ is:
$\Pr[g(X)\notin F(X)\text{ and }g(X^{\prime})\notin F(X^{\prime})\mid x^{-k}]=p_{g}(x^{-k})^{2}.$
Taking expectation over $x^{-k}$, we get:
$\Pr[g(X)\notin F(X)\text{ and }g(X^{\prime})\notin F(X^{\prime})]=\mathbb{E}_{x^{-k}}[p_{g}(x^{-k})^{2}]\geq\left(\mathbb{E}_{x^{-k}}[p_{g}(x^{-k})]\right)^{2}>\varepsilon^{2}.$
Under the event where both $g(X)\notin F(X)$ and $g(X^{\prime})\notin F(X^{\prime})$, we have $g(X)=h(x^{-k})\notin F(X)$ and $g(X^{\prime})=h(x^{-k})\notin F(X^{\prime})$. Therefore, $h(x^{-k})\notin F(X)\cap F(X^{\prime})$.
Since $g$ is constant in $x_{k}$, it assigns the same value $h(x^{-k})$ to both $X$ and $X^{\prime}$. No other function $g^{\prime}\in\mathcal{P}$ can assign a different value at $x^{-k}$, because all functions in $\mathcal{P}$ must be constant in the $k$-th dimension. Thus, there is no value $y$ such that $y\in F(X)\cap F(X^{\prime})$ that could correspond to $g^{\prime}(X)=y$ and $g^{\prime}\in\mathcal{P}$. Hence, $F(X)\cap F(X^{\prime})=\emptyset$, and the algorithm returns Reject.
Thus the algorithm Rejects with probability at least $\varepsilon^{2}$ and therefore the algorithm has detection probability $\rho(\varepsilon)=\varepsilon^{2}$ and satisfies Definition 4 for being a Proximity Oblivious Tester. ∎
B.2 Proximity Oblivious Tester for monotonic property
We prove that Algorithm 1 is indeed a POT for the property ${\mathcal{P}}$ of a function being monotonically decreasing in a given dimension $k$.
Proof.
Assume that $F$ accommodates $\mathcal{P}$. Then there exists a function $g\in\mathcal{P}$ such that for all $X\in\mathcal{X}$,
$g(X)\in F(X).$
Since $g$ is monotonically decreasing in $x_{k}$, for any fixed $x^{-k}\in\mathbb{R}^{d-1}$ and any $x_{k},x_{k}^{\prime}\in\mathbb{R}$ with $x_{k}\leq x_{k}^{\prime}$,
$g(x_{k},x^{-k})\geq g(x_{k}^{\prime},x^{-k}).$
In the algorithm, we sample $X_{1}=(x_{1}^{k},x^{-k})$ and $X_{2}=(x_{2}^{k},x^{-k})$, where $x^{-k}$ is fixed and $x_{1}^{k},x_{2}^{k}$ are independent samples from the marginal distribution of $x_{k}$. First consider the case that $x_{1}^{k}<x_{2}^{k}$. Since $g$ is monotonically decreasing, we have $g(X_{1})\geq g(X_{2})$. Because $g(X_{i})\in F(X_{i})$, it follows that
$\max F(X_{1})\geq g(X_{1})\geq g(X_{2})\geq\min F(X_{2}).$
Therefore, $\max F(X_{1})\geq\min F(X_{2})$, and the condition $\max F(X_{1})<\min F(X_{2})$ in step 3 is not satisfied. Now in the second case we have $x_{2}^{k}<x_{1}^{k}$. Similarly, $g(X_{2})\geq g(X_{1})$. Thus,
$\max F(X_{2})\geq g(X_{2})\geq g(X_{1})\geq\min F(X_{1}),$
implying $\max F(X_{2})\geq\min F(X_{1})$, so the condition in step 4 is not satisfied.
In both cases, the algorithm does not Reject and therefore Accepts. Thus, if $F$ accommodates $\mathcal{P}$, the algorithm always Accepts.
Now suppose $F$ is $\varepsilon$-far from $\mathcal{P}$. This means that for any $g\in\mathcal{P}$,
$\delta_{\mathcal{D}}(F,g)=\Pr_{X\sim\mathcal{D}}[g(X)\notin F(X)]>\varepsilon.$
For each $g\in\mathcal{P}$, define the set
$S_{g}=\{X\in\mathcal{X}:g(X)\notin F(X)\}.$
While $S_{g}$ may vary with $g$, we have $\mathcal{D}(S_{g})>\varepsilon$ for each $g\in\mathcal{P}$.
For a fixed $x^{-k}\in\mathbb{R}^{d-1}$, define
$p_{g}(x^{-k})=\Pr_{x_{k}\sim\mathcal{D}_{x_{k}|x^{-k}}}[g(x_{k},x^{-k})\notin F(x_{k},x^{-k})],$
where $\mathcal{D}_{x_{k}|x^{-k}}$ is the conditional distribution of $x_{k}$ given $x^{-k}$. Then,
$\mathbb{E}_{x^{-k}}[p_{g}(x^{-k})]=\Pr_{X\sim\mathcal{D}}[g(X)\notin F(X)]>\varepsilon.$
When we sample $X_{1}$ and $X_{2}$ with the same $x^{-k}$ and independent $x_{k}$ values, the probability that both $g(X_{1})\notin F(X_{1})$ and $g(X_{2})\notin F(X_{2})$ is
$\Pr[g(X_{1})\notin F(X_{1})\text{ and }g(X_{2})\notin F(X_{2})\mid x^{-k}]=p_{g}(x^{-k})^{2}.$
Taking expectation over $x^{-k}$, we get
$\Pr[g(X_{i})\notin F(X_{i})\text{ for }i=1,2]=\mathbb{E}_{x^{-k}}[p_{g}(x^{-k})^{2}]\geq\left(\mathbb{E}_{x^{-k}}[p_{g}(x^{-k})]\right)^{2}>\varepsilon^{2},$
Under the event where both $g(X_{1})\notin F(X_{1})$ and $g(X_{2})\notin F(X_{2})$, consider the following cases. In the first case, $x_{1}^{k}<x_{2}^{k}$.Since $g$ is monotonically decreasing, $g(X_{1})\geq g(X_{2})$. However, $g(X_{1})\notin F(X_{1})$ and $g(X_{2})\notin F(X_{2})$. To prevent any $g^{\prime}\in\mathcal{P}$ from satisfying $g^{\prime}(X_{i})\in F(X_{i})$, it must be that $\max F(X_{1})<\min F(X_{2})$. If $\max F(X_{1})\geq\min F(X_{2})$, a function $g^{\prime}$ could exist with $g^{\prime}(X_{i})\in F(X_{i})$ and satisfying monotonicity. In the second case, $x_{2}^{k}<x_{1}^{k}$, similar reasoning leads to $\max F(X_{2})<\min F(X_{1})$. Therefore, when both $g(X_{i})\notin F(X_{i})$ and the appropriate ordering of $x_{k}$ holds, the algorithm Rejects.
Since the probability of both $g(X_{i})\notin F(X_{i})$ occurring is at least $\varepsilon^{2}$, and the conditions for rejection are met under these events, the algorithm Rejects with probability at least $\varepsilon^{2}$.
Thus, Algorithm 1 satisfies the conditions of a Proximity-Oblivious Tester for the property $\mathcal{P}$ of monotonically decreasing in a given dimension $k$.
B.3 Proximity Oblivious Tester for concave property
Proof: Assume that $F$ accommodates $\mathcal{P}$; that is, there exists a function $g\in\mathcal{P}$ such that for all $X\in\mathcal{X}$, $g(X)\in F(X).$ Since $g$ is concave in the $k$-th dimension, for any $X_{\text{left}},X_{\text{mid}},X_{\text{right}}$ constructed as in the algorithm, the following inequality holds:
$g(X_{\text{mid}})\leq\alpha g(X_{\text{left}})+(1-\alpha)g(X_{\text{right}}),$
where
$\alpha=\frac{X_{\text{right}}[k]-X_{\text{mid}}[k]}{X_{\text{right}}[k]-X_{\text{left}}[k]}.$
Because $g(X_{i})\in F(X_{i})$ for $i\in\{\text{left},\text{mid},\text{right}\}$, we have:
$\min F(X_{\text{mid}})$ $\leq g(X_{\text{mid}}),$
$g(X_{\text{left}})$ $\leq\max F(X_{\text{left}}),$
$g(X_{\text{right}})$ $\leq\max F(X_{\text{right}}).$
Substituting these into the concavity inequality, we obtain:
$\min F(X_{\text{mid}})\leq\alpha\max F(X_{\text{left}})+(1-\alpha)\max F(X_{\text{right}}).$
Therefore, the condition in line 7 of the algorithm is not satisfied (i.e., the inequality for rejection does not hold), and the algorithm returns Accept.
Now suppose that $F$ is $\varepsilon$-far from $\mathcal{P}$. This means that for any $g\in\mathcal{P}$,
$\delta_{\mathcal{D}}(F,g)=\Pr_{X\sim\mathcal{D}}[g(X)\notin F(X)]>\varepsilon.$
For each $g\in\mathcal{P}$, define the set:
$S_{g}=\{X\in\mathcal{X}:g(X)\notin F(X)\}.$
While $S_{g}$ may vary with $g$, we have $\mathcal{D}(S_{g})>\varepsilon$ for each $g\in\mathcal{P}$. The probability that $g(X_{\text{mid}})\notin F(X_{\text{mid}})$ is greater than $\varepsilon$, and similarly, the probabilities that $g(X_{\text{left}})\notin F(X_{\text{left}})$ and $g(X_{\text{right}})\notin F(X_{\text{right}})$ are each greater than $\varepsilon$. Thus, by independence of the samples, the joint probability that all three events occur is at least $\varepsilon^{3}$:
$\Pr\big{[}g(X_{i})\notin F(X_{i})\text{ for all }i\in\{\text{left},\text{mid},\text{right}\}\big{]}\geq\varepsilon^{3}.$
Under the event where $g(X_{i})\notin F(X_{i})$ for all $i\in\{\text{left},\text{mid},\text{right}\}$, we will show that no concave function $g^{\prime}\in\mathcal{P}$ can satisfy $g^{\prime}(X_{i})\in F(X_{i})$ for all $i$.
Assume for contradiction that there exists a concave function $g^{\prime}\in\mathcal{P}$ such that
$g^{\prime}(X_{i})\in F(X_{i})\quad\text{for all }i\in\{\text{left},\text{mid},\text{right}\}.$
Since $g^{\prime}$ is concave in the $k$-th dimension, it must satisfy:
$g^{\prime}(X_{\text{mid}})\leq\alpha g^{\prime}(X_{\text{left}})+(1-\alpha)g^{\prime}(X_{\text{right}}).$
Using the bounds from $F(X_{i})$:
$g^{\prime}(X_{\text{mid}})$ $\geq\min F(X_{\text{mid}}),$
$g^{\prime}(X_{\text{left}})$ $\leq\max F(X_{\text{left}}),$
$g^{\prime}(X_{\text{right}})$ $\leq\max F(X_{\text{right}}).$
Substituting these into the concavity inequality:
$\min F(X_{\text{mid}})\leq\alpha\max F(X_{\text{left}})+(1-\alpha)\max F(X_{\text{right}}).$
However, the algorithm’s condition in line 7 specifies that:
$\min F(X_{\text{mid}})>\alpha\max F(X_{\text{left}})+(1-\alpha)\max F(X_{\text{right}}).$
This is a contradiction. Therefore, our assumption is false; no concave function $g^{\prime}\in\mathcal{P}$ exists that satisfies $g^{\prime}(X_{i})\in F(X_{i})$ for all $i$. Since under this event (which occurs with probability at least $\varepsilon^{3}$) no concave function $g^{\prime}\in\mathcal{P}$ can fit within $F(X_{i})$ at the points $X_{\text{left}},X_{\text{mid}},X_{\text{right}}$, the algorithm correctly Rejects.
Therefore, the algorithm Rejects with probability at least $\varepsilon^{3}$, so the algorithm satisfies the conditions of a Proximity-Oblivious Tester for the property $\mathcal{P}$ of concavity in the $k$-th dimension.
Appendix C Proof of multi-lambda conformal risk control
We emphasize that this claim follows directly from the original proof and result of conformal risk control [Angelopoulos et al., 2024] and we present it here mainly for completeness given our use of it in this multi-dimensional case.